# Design of Bayesian A/B Tests Controlling False Discovery Rates and Power

### v3

@@ 1 Introduction
The rise of e-commerce is indisputable – accounting for 20% of global commerce in 2024 and expected to reach 23% by 2028 (Cramer-Flood, 2024). To stay competitive amid the changing commercial landscape, it is important for online businesses and brick-and-mortar businessess with online components to optimize their products and services. Technology-based companies such as Airbnb, Facebook, LinkedIn, Microsoft, Netflix, Spotify, and Uber all rely on experimentation to optimize how they advertise their services, acquire and retain customers, and develop products (Thomke, 2020; Luca and Bazerman, 2021). This optimization process is carried out by testing different versions of products and services through online controlled experiments. These experiments are often informally called A/B tests, which is an apt moniker for comparisons involving two versions of an online experience: a control version (A) and one treatment version (B). Although the design of online controlled experiments can be substantially more complicated (see e.g., Kohavi et al. (2020)), the majority of these experiments are traditional A/B tests, so that is the focus of this paper.
Several reviews of A/B testing, each with a distinct focus, have recently been conducted. Gupta et al. (2019) discussed A/B testing methods for industry professionals with an emphasis on operational challenges. Stevens and Hagar (2022) highlighted the use of Bayesian A/B tests for practitioners. Bojinov and Gupta (2022) detailed best practices for business leaders to scale A/B testing. Larsen et al. (2024) overviewed statistical challenges associated with A/B testing for academic statisticians to promote collaboration between academia and industry. Readers are directed to these reviews and the references therein for additional background on A/B testing methods.
A/B tests compare two versions of an online experience using a collection of metrics, where each metric numerically summarizes a different aspect of how users interact with a product or service. In A/B testing, a vast collection of metrics is typically monitored (Larsen et al., 2024; Bajpai et al., 2025). Debugging metrics help diagnose implementation problems, guardrail metrics safeguard against negative business impacts, and a limited set of metrics defines the overall evaluation criterion (OEC) used for decision making. For various A/B tests, a weighted combination of the metrics in this limited set defines a (composite) scalar OEC (Roy, 2001). While using one scalar OEC simplifies decision making, it can be difficult to adequately summarize the business goals for an A/B test into a single metric.
In practice, many companies instead examine a multi-metric OEC comprised of several important scalar metrics. Decision-making procedures that rely on multiple metrics are sometimes called balanced scorecard approaches (Kaplan and Norton, 1996). Kohavi et al. (2020) provided a rough rule of thumb to limit the number of key metrics to at most $K=5$ to account for the multiple comparisons problem (Hochberg and Tamhane, 2009). The multiple comparisons problem addresses the growing risk of making erroneous inferences as the number of statistical comparisons increases. For each metric $k=1,\dots,K$ considered, complementary hypotheses $H_{0,k}$ and $H_{1,k}$ are compared. A discovery is made with respect to the $k^{\text{th}}$ metric when the data support $H_{1,k}$. The rule of thumb with five metrics for A/B tests was provided based on guidance for controlling the family-wise error rate (FWER). The FWER is the probability of making at least one type I error (i.e., incorrectly concluding at least one $H_{1,k}$ is true) across all $K$ hypotheses examined.
@@
However, strict control of the FWER often results in a loss of power to detect differences between online experiences. Insufficient power in A/B tests is already a common and well-documented problem (Deng et al., 2013; Larsen et al., 2024; Deng et al., 2024), so any design choices that further decrease power should be avoided. As such, we consider controlling the false discovery rate (FDR): the expected proportion of discoveries that are false across a group of hypothesis tests (Benjamini and Hochberg, 1995). Whereas the FWER conservatively bounds the probability of making even one false discovery, the FDR balances the prevalence of false and true positives across multiple comparisons. Recent contributions have considered the FDR for A/B tests in a post hoc manner (Kohavi and Chen, 2024). In particular, Berman et al. (2018) and Berman and Van den Bulte (2022) estimated that roughly 20% of A/B tests conducted at a 5% significance level result in false discoveries. The development of more principled methods for designing A/B tests with FDR control is therefore crucial to enhance the power of these tests.
The most popular methods used to control the FDR are based on frequentist $p$-values (Benjamini and Hochberg, 1995; Storey, 2002). These standard inference methods require that all hypotheses are independent, which may be an unrealistic assumption for many A/B tests. While analytical methods to control the FDR under arbitrary dependence assumptions have been developed (Benjamini and Yekutieli, 2001), simulation-based procedures are recommended for optimal control of the FDR for a specific dependence structure (Yekutieli and Benjamini, 1999). Under independence assumptions, Benjamini and Hochberg (1995) demonstrated that hypothesis testing with the FDR yielded higher average power than testing methods that control the FWER. Average power is the expected proportion of the true $\{H_{1,k}\}_{k=1}^{K}$ that are correctly supported by the data. From a design perspective, methods to select sample sizes that attain a desired average power have been developed in frequentist settings with independent hypothesis tests (Jung, 2005; Pounds and Cheng, 2005).
Bayesian A/B tests are becoming increasingly popular in part due to the statistical community’s concern with the misuse of $p$-values (Wasserstein and Lazar, 2016; Wasserstein et al., 2019) and because they offer several advantages relative to frequentist methods. For instance, Bayesian methods easily account for prior information, which is abundant in online settings. Bayesian methods, particularly posterior probabilities, are also more readily understood by practitioners who conduct A/B tests. A/B testing methods with posterior probabilities have been proposed in several settings (see e.g., Deng (2015); Stevens and Hagar (2022); Deng et al. (2024)). For the $k^{\text{th}}$ metric in a group of $K$ comparisons, the observed data provide sufficient evidence to support $H_{1,k}$ if the posterior probability $Pr(H_{1,k}~{}|~{}data)$ is greater than or equal to a decision threshold $\gamma\in[0.5,1)$. Given some data generation process, a sample size $n$ could be selected to attain a value for average power as defined by Benjamini and Hochberg (1995) and described above. In that event, the selected $n$ should ensure the average probability that $Pr(H_{1,k}~{}|~{}data)\geq\gamma$ across all true hypotheses in $\{H_{1,k}\}_{k=1}^{K}$ is at least $1-\beta$. The decision threshold $\gamma$ is chosen to bound the FDR by $q\in(0,1)$. Regulatory agencies require strict control of operating characteristics in clinical settings (FDA, 2019), but Bayesian A/B tests may also be designed using frequentist operating characteristics (Stevens and Hagar, 2022; Deng et al., 2024). Although not the focus of this paper, alternative Bayesian approaches to select sample sizes under FDR control that involve decision theory have been proposed (Müller et al., 2004, 2007).
@@
To support the design of A/B tests with non-simplistic statistical models, sample sizes and decision thresholds that control the operating characteristics of average power and the FDR can be found using intensive simulation (Wang and Gelfand, 2002). In general, one must simulate many samples of size $n$ according to some assumed data generation process to estimate the joint sampling distribution of $\{Pr(H_{1,k}~{}|~{}data)\}_{k=1}^{K}$. This estimate of the sampling distribution of posterior probabilities determines whether the investigated sample size and decision threshold satisfy criteria for the operating characteristics. This process is repeated until a suitable $(n,\gamma)$ combination is found. This computational burden is compounded over all data generation processes practitioners want to consider. An economical framework to determine optimal sample sizes and decision thresholds that satisfy criteria for the FDR and average power would expedite the design of Bayesian A/B tests with multiple metrics, making them more accessible to practitioners.
@@
Several strategies have recently been developed to accelerate study design based on sampling distributions of posterior probabilities. Approaches that efficiently fit beta distributions to sampling distributions of posterior probabilities have been developed in clinical settings (Golchi, 2022; Golchi and Willard, 2024). Hagar and Stevens (2024) prioritized exploring segments of the sampling distribution of posterior probabilities that are near the decision threshold for a single hypothesis test. Hagar and Stevens (2025) put forward a method to assess operating characteristics throughout the sample size space for a single test using estimates of the sampling distribution at only two sample sizes. Hagar and Golchi (2025) extended this method to accommodate clustered data and several hypotheses in clinical settings without considering the multiple comparisons problem. In this work, we build upon the methods from Hagar and Stevens (2025) and Hagar and Golchi (2025) to more efficiently design A/B tests involving a host of metrics controlling the FDR and average power. Because none of the aforementioned strategies were proposed with A/B tests or inflated type I error rates in mind, there is a real need for the methodology proposed in this paper.
@@
The remainder of this article is structured as follows. We introduce background information and notation in Section 2. In Section 3, we prove theoretical results about a proxy to the joint sampling distribution of posterior probabilities. We develop a method in Section 4 that adapts these results to determine which $(n,\gamma)$ combination minimizes the sample size $n$ while satisfying criteria for average power and the FDR. This method efficiently considers a range of sample sizes using estimates of the sampling distributions of posterior probabilities at only two sample sizes. We also describe a bootstrap procedure to quantify the impact of simulation variability on the sample size recommendation. In Section 5, we illustrate the use of our methodology with an example from a real A/B test with multiple binary metrics that have a complicated dependence structure. In Section 6, we repurpose the posterior probabilities used to find the optimal $(n,\gamma)$ combination to create contour plots that facilitate the investigation of various $n$ and $\gamma$ values. We conclude with a summary and discuss extensions to this work in Section 7.
@@ 2 Preliminaries
Our design framework represents data from a to-be-observed sample of $n$ observations as $\boldsymbol{Y}^{{}_{(n)}}=\{\boldsymbol{Y}_{i}\}_{i=1}^{n}$. We presume there is one outcome for each of the $K$ metrics, so each observation $\boldsymbol{Y}_{i}$ has dimension $K$. The observed data are denoted by $\boldsymbol{y}^{{}_{(n)}}$. For A/B tests, $\boldsymbol{Y}^{{}_{(n)}}$ consists of $n=n_{A}+n_{B}$ observations, where the first $n_{A}$ observations are from group A and the final $n_{B}$ observations are from group B. We consider fixed treatment allocation such that $n_{A}=\lfloor cn_{B}\rceil$ for some $c>0$, but this constant $c$ is not incorporated into $\boldsymbol{Y}^{{}_{(n)}}$. We assume that each observation in $\boldsymbol{Y}^{{}_{(n)}}$ is generated independently according to the model $f(\boldsymbol{y};\boldsymbol{\eta}_{A})$ or $f(\boldsymbol{y};\boldsymbol{\eta}_{B})$, depending on whether the observation came from group A or B. We note that $\boldsymbol{\eta}_{j}$ represents a vector of parameters for group $j=A,B$, and we collectively denote the parameters for the two groups by $\boldsymbol{\eta}=(\boldsymbol{\eta}_{A},\boldsymbol{\eta}_{B})$. This paper focuses on A/B tests that do not adjust for additional covariates $\{\boldsymbol{X}_{i}\}_{i=1}^{n}$, which would not be available when users interact anonymously with an online platform (Sim et al., 2024). However, Hagar and Stevens (2025) considered designs with additional covariates, so our extensions to their method could also be used to design A/B tests that involve treatment effect heterogeneity (Taddy et al., 2016).
@@
The targets of inference $\boldsymbol{\theta}$ for the group of metrics are specified as a function $g(\cdot)$ of the model parameters: $\boldsymbol{\theta}=g(\boldsymbol{\eta})$. The components in $\boldsymbol{\theta}=\{\theta_{k}\}_{k=1}^{K}$ can be expressed as $\theta_{k}=g_{k}(\boldsymbol{\eta})$. For the $k^{\text{th}}$ metric, we consider interval hypotheses of the form $H_{1,k}:\theta_{k}\in(\delta_{L,k},\delta_{U,k})$, where $-\infty\leq\delta_{L,k}<\delta_{U,k}\leq\infty$. The corresponding complementary hypothesis is $H_{0,k}:\theta_{k}\notin(\delta_{L,k},\delta_{U,k})$. General notation for the interval $(\delta_{L,k},\delta_{U,k})$ is used to accommodate a broad suite of hypothesis tests (Stevens and Hagar, 2022). We assume larger $\theta_{k}$ values are preferred to introduce several such hypotheses. In that case, the interval $(\delta_{L,k},\delta_{U,k})=(0,\infty)$ facilitates one-sided hypothesis tests based on superiority. The interval endpoint $\delta_{L,k}$ might instead take a negative (positive) value to assess hypotheses based on noninferiority (practical superiority). For an equivalence test, both interval endpoints of $(\delta_{L,k},\delta_{U,k})$ would take finite values. We jointly refer to the interval endpoints across all hypotheses as $\boldsymbol{\delta}_{L}=\{\delta_{L,k}\}_{k=1}^{K}$ and $\boldsymbol{\delta}_{U}=\{\delta_{U,k}\}_{k=1}^{K}$.
@@
Algorithm 1 details a simulation-driven procedure to classify hypotheses as true or false for a particular $(n,\gamma)$ combination and set of design inputs. These classifications are based on posterior probabilities, and they are used to estimate the FDR and average power. We now discuss the design inputs for this algorithm that were not described above. The prior $p(\boldsymbol{\eta})$ is used to define the posterior of $\boldsymbol{\theta}$. We must also specify the number of simulation repetitions $m$. The final design input that we must specify is $\Psi$, a probability model that characterizes how $\boldsymbol{\eta}$ values are drawn in each simulation repetition $r=1,\dots,m$. The model $\Psi$ defines scenarios with effect sizes that align with the experiment’s objectives. It differs from the prior $p(\boldsymbol{\eta})$ that characterizes all available pre-experimental knowledge regardless of whether it aligns with the goals of the A/B tests. Note that the FDR is trivially 0 if all hypotheses $\{H_{1,k}\}_{k=1}^{K}$ are true, and average power cannot be defined when all $\{H_{1,k}\}_{k=1}^{K}$ are false. When designing A/B tests using the FDR and average power, all $\boldsymbol{\eta}_{r}\sim\Psi$ must therefore be such that some – but not all – of $\{H_{1,k}\}_{k=1}^{K}$ are false. This assumption about $\{H_{1,k}\}_{k=1}^{K}$ is used for design purposes; there are no restrictions on how many hypotheses are false when using posterior probabilities to assess a collection of A/B tests. As illustrated in Section 5, the model $\Psi$ can incorporate uncertainty about how many and which hypotheses are false.
@@
1:procedure Classify($f(\cdot)$, $g(\cdot)$, $\boldsymbol{\delta}_{L}$, $\boldsymbol{\delta}_{U}$, $n$, $c$, $\gamma$, $K$, $p(\boldsymbol{\eta})$, $m$, $\Psi$)
@@
9:                  if $\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}}_{r})\geq\gamma$ then
@@
13:                  if $\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}}_{r})\geq\gamma$ then
@@
$\dfrac{1}{m}\sum_{r=1}^{m}\dfrac{v_{r}}{\max\{v_{r}+s_{r},1\}}~{}~{}~{}\text{and}~{}~{}~{}\dfrac{1}{m}\sum_{r=1}^{m}\dfrac{s_{r}}{t_{r}},$ (1)
@@
where $v_{r}$, $s_{r}$, and $t_{r}$ are the number of false discoveries, true discoveries, and true hypotheses for simulation repetition $r$ defined in Algorithm 1. The denominator of the FDR in (1) avoids division by 0, aligning with the formal definition from Benjamini and Hochberg (1995). Design with modified versions of the FDR, such as the positive FDR (Storey, 2003), could also be considered by adjusting the equations for the estimates in (1). Based on Algorithm 1, an $(n,\gamma)$ combination is suitable if the FDR estimate in (1) is at most $q$ and the average power estimate in (1) is at least $1-\beta$.
@@
We now separately consider how the marginal sampling distribution of $Pr(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}})$ for each hypothesis $H_{1,k}$ behaves with respect to $n$. We first suppose the model $\Psi$ is defined such that $H_{1,k}$ is true for some $\boldsymbol{\eta}_{r}$ values and false for others. In this case, the marginal sampling distribution of $Pr(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}})$ has two components. The first component is based on $Pr(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}}_{r})$ for $\boldsymbol{\eta}_{r}\sim\Psi$ such that $H_{1,k}$ is true. As $n\rightarrow\infty$, this component of the marginal sampling distribution generally converges to a point mass at 1 (van der Vaart, 1998). The second component is based on $Pr(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}}_{r})$ for $\boldsymbol{\eta}_{r}\sim\Psi$ such that $H_{1,k}$ is false. When all those $\boldsymbol{\eta}_{r}$ values are such that $\theta_{r,k}=g_{k}(\boldsymbol{\eta}_{r})$ equals $\delta_{L,k}$ or $\delta_{U,k}$, this component of the marginal sampling distribution for $H_{1,k}$ converges to the standard uniform distribution as $n\rightarrow\infty$ under weak conditions (Bernardo and Smith, 2009). If $\Psi$ is defined such that $H_{1,k}$ is true or false for all $\boldsymbol{\eta}_{r}$ values, the marginal sampling distribution of $Pr(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}})$ contains only one of these two components.
@@
In all cases discussed above, the optimal value for the decision threshold $\gamma$ is the smallest probability that bounds the estimated FDR in (1) by $q$. As $n$ increases, so do the posterior probabilities in the components of the marginal sampling distributions where $H_{1,k}$ is true, and it becomes easier to distinguish between the true and false hypotheses. The optimal $\gamma$ value typically decreases alongside $n$; however, this decreasing trend is attenuated as $n\rightarrow\infty$ since the components of the marginal sampling distributions of posterior probabilities where $H_{1,k}$ is true converge to 1. Even as the optimal $\gamma$ changes, average power increases to 1 as $n$ increases. The optimal choices for $n$ and $\gamma$ are therefore intrinsically linked when designing A/B tests based on the FDR and average power.
@@
These choices for $n$ and $\gamma$ are informed by exploring the joint sampling distribution of posterior probabilities under $\Psi$ to estimate the operating characteristics in (1). While we can examine multiple $\gamma$ values using the same estimate of the sampling distribution for a given sample size $n$, the procedure in Algorithm 1 requires independent implementation for each $n$ value we consider. This process is often computationally intensive, but we could reduce the computational burden by repurposing the results from Algorithm 1 for previously considered sample sizes to estimate operating characteristics for new $n$ values. This process would allow us to explore new $(n,\gamma)$ combinations without conducting additional simulations. We propose such a method for the design of A/B tests in this paper and begin its development in Section 3.

### v4

@@ 1 Introduction
In many contexts involving business and the social sciences, statisticians aim to assess the impact of an intervention on multiple outcomes. One prevalent context is that of online controlled experiments run by technology-based companies to improve their products and services (Thomke, 2020; Luca and Bazerman, 2021; Bajpai et al., 2025). These experiments are often called A/B tests, which is an apt moniker for comparisons involving two versions of an online experience: a control version (A) and a treatment version (B). A/B tests typically monitor many metrics that summarize different aspects of how users interact with an online experience. A limited set of these metrics defines the overall evaluation criterion (OEC) used for decision making (Kohavi et al., 2020). Combining this set of metrics into one composite OEC simplifies decision making (Roy, 2001), but adequately summarizing the business goals for an A/B test with a single metric can present challenges. Many companies instead examine a multi-metric OEC comprised of several key scalar metrics, and so the multiple comparisons problem (Hochberg and Tamhane, 2009) must be accounted for during A/B test design. For further background on A/B testing methods, see Larsen et al. (2024).
@@
For each metric $k=1,\dots,K$, A/B tests compare complementary hypotheses $H_{0,k}$ and $H_{1,k}$. A discovery is made with respect to the $k^{\text{th}}$ metric when the data support $H_{1,k}$. A/B tests with multiple metrics are often designed to control the family-wise error rate (FWER). However, strict control of the FWER often results in a loss of power. Since insufficient power in A/B tests is a common problem (Deng et al., 2013; Larsen et al., 2024; Deng et al., 2024), any design choices that further decrease power should be avoided. Thus, we control the false discovery rate (FDR): the expected proportion of discoveries that are false across a group of hypothesis tests (Benjamini and Hochberg, 1995). Whereas the FWER conservatively bounds the probability of making even one false discovery (i.e., incorrectly concluding that one $H_{1,k}$ is true), the FDR balances the prevalence of false and true positives across multiple comparisons. The development of design methods for A/B tests with FDR control is therefore crucial to enhance the average power of these tests. Average power is the expected proportion of the true $\{H_{1,k}\}_{k=1}^{K}$ that are correctly supported by the data.
The most popular methods used to control the FDR are based on frequentist $p$-values (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001; Storey, 2002). Given the rising prominence of Bayesian approaches to A/B testing (see e.g., Deng (2015); Stevens and Hagar (2022); Deng et al. (2024)), this paper focuses on testing procedures facilitated via posterior probabilities. For the $k^{\text{th}}$ metric, the observed data provide evidence to support $H_{1,k}$ if the posterior probability $Pr(H_{1,k}~{}|~{}data)$ is greater than or equal to a decision threshold $\gamma_{k}\in[0.5,1)$. Given some data generation process, a sample size $n$ could be selected to ensure the average power to observe $Pr(H_{1,k}~{}|~{}data)\geq\gamma_{k}$ across all true hypotheses in $\{H_{1,k}\}_{k=1}^{K}$ is at least $1-\beta$. The decision thresholds $\boldsymbol{\gamma}=\{\gamma_{k}\}_{k=1}^{K}$ are chosen to bound the FDR by $q\in(0,1)$. The FDR control provided by this framework with posterior probabilities is compared with the FDR control prompted by popular frequentist methods in Section 2.
@@
To support the design of A/B tests with non-simplistic statistical models, sample sizes and optimal decision thresholds that control the operating characteristics of average power and the FDR can be found using intensive simulation (Wang and Gelfand, 2002). In general, one must simulate many samples of size $n$ according to some assumed data generation process to estimate the joint sampling distribution of $\{Pr(H_{1,k}~{}|~{}data)\}_{k=1}^{K}$. This estimate of the sampling distribution of posterior probabilities determines whether the investigated sample size and decision thresholds satisfy criteria for the operating characteristics. This computationally intensive process is repeated until a suitable $(n,\boldsymbol{\gamma})$ combination is found. An economical framework to determine optimal sample sizes and decision thresholds that satisfy criteria for the FDR and average power would expedite the design of Bayesian A/B tests with multiple metrics, making them more accessible to practitioners.
@@
For a single test, Hagar and Stevens (2025) recently put forward an economical method to assess operating characteristics throughout the sample size space using estimates of the sampling distribution of posterior probabilities at only two sample sizes. In this work, we build upon their approach to more efficiently design A/B tests involving a host of metrics controlling the FDR and average power. To assess the operating characteristics of multiple tests, we must account for the dependence structure in the joint sampling distribution of posterior probabilities across hypotheses. The method from Hagar and Stevens (2025) was predicated on a theoretical proxy to the (univariate) sampling distribution of posterior probabilities for a single hypothesis. As such, new theory must be developed to accommodate arbitrary dependence structures in the joint sampling distribution.
@@
The remainder of this article is structured as follows. We introduce our framework for inference and the criteria under which the FDR and average power are controlled in Section 2. In Section 3, we prove novel theoretical results about a proxy to the joint sampling distribution of posterior probabilities. We develop a method in Section 4 that adapts these results to determine which $(n,\boldsymbol{\gamma})$ combination minimizes the sample size $n$ while satisfying criteria for average power and the FDR. This method efficiently considers a range of sample sizes using estimates of the sampling distributions of posterior probabilities at only two values of $n$. In Section 5, we illustrate the use of our methodology with an example from a real A/B test with multiple binary metrics that have a complicated dependence structure. We conclude with a summary and discuss extensions to this work in Section 6.
@@ 2.1 Problem Set-Up
Our design framework represents data from a future sample of size $n$ as $\boldsymbol{Y}^{{}_{(n)}}=\{\boldsymbol{Y}_{i}\}_{i=1}^{n}$. We presume there is one outcome for each of the $K$ metrics, so each observation $\boldsymbol{Y}_{i}$ has dimension $K$. The observed data are denoted by $\boldsymbol{y}^{{}_{(n)}}$. For A/B tests, $\boldsymbol{Y}^{{}_{(n)}}$ consists of $n_{j}$ observations from group $j=A,B$ such that $n=n_{A}+n_{B}$. We consider fixed treatment allocation where $n_{A}=\lfloor cn_{B}\rceil$ for some $c>0$, but this constant $c$ is not incorporated into $\boldsymbol{Y}^{{}_{(n)}}$. We assume that each observation in $\boldsymbol{Y}^{{}_{(n)}}$ is generated independently via the model $f(\boldsymbol{y};\boldsymbol{\eta}_{A})$ or $f(\boldsymbol{y};\boldsymbol{\eta}_{B})$, depending on its group assignment. We note that $\boldsymbol{\eta}_{j}$ represents a vector of parameters for group $j=A,B$, and we collectively denote the parameters for both groups by $\boldsymbol{\eta}=(\boldsymbol{\eta}_{A},\boldsymbol{\eta}_{B})$. Since Hagar and Stevens (2025) considered designs with additional covariates $\{\boldsymbol{X}_{i}\}_{i=1}^{n}$, our extensions to their method could also be used to design A/B tests that involve treatment effect heterogeneity (Larsen et al., 2024).
@@
The targets of inference $\boldsymbol{\theta}$ for the host of metrics are specified as a function $g(\cdot)$ of the model parameters: $\boldsymbol{\theta}=g(\boldsymbol{\eta})$. The components in $\boldsymbol{\theta}=\{\theta_{k}\}_{k=1}^{K}$ can be expressed as $\theta_{k}=g_{k}(\boldsymbol{\eta})$. For the $k^{\text{th}}$ metric, we consider interval hypotheses of the form $H_{1,k}:\theta_{k}\in(\delta_{L,k},\delta_{U,k})$, where $-\infty\leq\delta_{L,k}<\delta_{U,k}\leq\infty$. The corresponding complementary hypothesis is $H_{0,k}:\theta_{k}\notin(\delta_{L,k},\delta_{U,k})$. This general notation for the interval $(\delta_{L,k},\delta_{U,k})$ accommodates a broad suite of hypothesis tests based on superiority and practical equivalence (Stevens and Hagar, 2022). We jointly refer to the interval endpoints across all hypotheses as $\boldsymbol{\delta}_{L}=\{\delta_{L,k}\}_{k=1}^{K}$ and $\boldsymbol{\delta}_{U}=\{\delta_{U,k}\}_{k=1}^{K}$.
@@
Algorithm 1 details a simulation-driven procedure to classify hypotheses as true or false for a particular $(n,\boldsymbol{\gamma})$ combination and set of design inputs. These classifications based on posterior probabilities are used to estimate the FDR and average power. We now discuss several design inputs for this algorithm. The prior $p(\boldsymbol{\eta})$ is used to define the posterior of $\boldsymbol{\theta}$. We also specify $\Psi$, a probability model that characterizes how $\boldsymbol{\eta}$ values are drawn in each simulation repetition $r=1,\dots,m$. The model $\Psi$ defines scenarios with effect sizes that align with the experiment’s objectives. It differs from the prior $p(\boldsymbol{\eta})$ that characterizes all available pre-experimental knowledge regardless of whether it aligns with the goals of the A/B tests. As illustrated in Section 5, defining a model $\Psi$ that is distinct from $p(\boldsymbol{\eta})$ allows practitioners to incorporate uncertainty about how many and which hypotheses are false in a straightforward manner. Note that the FDR is trivially 0 if all hypotheses $\{H_{1,k}\}_{k=1}^{K}$ are true, and average power is trivially 0 when all $\{H_{1,k}\}_{k=1}^{K}$ are false. In this paper, we therefore focus on data generation processes where each $\boldsymbol{\eta}_{r}\sim\Psi$ is such that some – but not all – of $\{H_{1,k}\}_{k=1}^{K}$ are false. We use this assumption about $\{H_{1,k}\}_{k=1}^{K}$ for design purposes to ensure the FDR and average power calculations are non-trivial in all simulation repetitions; there are no restrictions on how many hypotheses are false when using posterior probabilities to assess a collection of A/B tests. We note that one could also use our design framework with a model $\Psi$ such that all $\{H_{1,k}\}_{k=1}^{K}$ are true or false in some simulation repetitions $r=1,\dots,m$.
@@
1:procedure Classify($f(\cdot)$, $g(\cdot)$, $\boldsymbol{\delta}_{L}$, $\boldsymbol{\delta}_{U}$, $n$, $c$, $\boldsymbol{\gamma}$, $K$, $p(\boldsymbol{\eta})$, $m$, $\Psi$)
@@
9:                  if $\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}}_{r})\geq\gamma_{k}$ then
@@
13:                  if $\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}}_{r})\geq\gamma_{k}$ then
@@
$\dfrac{1}{m}\sum_{r=1}^{m}\dfrac{v_{r}}{\max\{v_{r}+s_{r},1\}}~{}~{}~{}\text{and}~{}~{}~{}\dfrac{1}{m}\sum_{r=1}^{m}\dfrac{s_{r}}{\max\{t_{r},1\}},$ (1)
@@
where $v_{r}$, $s_{r}$, and $t_{r}$ are the number of false discoveries, true discoveries, and true hypotheses for simulation repetition $r$ defined in Algorithm 1. The denominators in (1) avoid division by 0, aligning with the definition of the FDR from Benjamini and Hochberg (1995). Based on Algorithm 1, an $(n,\boldsymbol{\gamma})$ combination is suitable if the FDR estimate in (1) is at most $q$ and the average power estimate in (1) is at least $1-\beta$.
@@
We now separately consider how the marginal sampling distribution of $Pr(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}})$ for each hypothesis $H_{1,k}$ behaves with respect to $n$. The marginal sampling distribution of $Pr(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}})$ has at least one of the following two components. The first component is based on $Pr(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}}_{r})$ for $\boldsymbol{\eta}_{r}\sim\Psi$ such that $H_{1,k}$ is true. As $n\rightarrow\infty$, this component of the marginal sampling distribution generally converges to a point mass at 1 (van der Vaart, 1998). The second component is based on $Pr(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n)}}_{r})$ for $\boldsymbol{\eta}_{r}\sim\Psi$ such that $H_{1,k}$ is false. When all those $\boldsymbol{\eta}_{r}$ values are such that $\theta_{r,k}=g_{k}(\boldsymbol{\eta}_{r})$ equals $\delta_{L,k}$ or $\delta_{U,k}$, this component of the marginal sampling distribution for $H_{1,k}$ converges to the standard uniform distribution as $n\rightarrow\infty$ under weak conditions (Bernardo and Smith, 2009).
@@
The optimal decision thresholds $\boldsymbol{\gamma}$ that bound the estimated FDR in (1) by $q$ can be found using a defined optimization scheme. We illustrate the use of several constrained optimize schemes that select $\boldsymbol{\gamma}$ to maximize average power in Section 5. As $n$ increases, so do the posterior probabilities in the components of the marginal sampling distributions where $H_{1,k}$ is true, and it becomes easier to distinguish between the true and false hypotheses. The optimal $\boldsymbol{\gamma}$ values for a given optimization scheme typically decrease alongside $n$; however, this decreasing trend is attenuated as $n\rightarrow\infty$ since the components of the marginal sampling distributions of posterior probabilities where $H_{1,k}$ is true converge to 1. Even as the optimal $\boldsymbol{\gamma}$ values change, average power increases to 1 as $n$ increases. The optimal choices for $n$ and $\boldsymbol{\gamma}$ are therefore intrinsically linked when designing A/B tests based on the FDR and average power.
@@
These choices for $n$ and $\boldsymbol{\gamma}$ are informed by exploring the joint sampling distribution of posterior probabilities under $\Psi$ to estimate the operating characteristics in (1). While we can examine multiple $\boldsymbol{\gamma}$ values using the same estimate of the sampling distribution for a given sample size $n$, the procedure in Algorithm 1 requires independent implementation for each $n$ value we consider. This process is often computationally intensive, but we could reduce the computational burden by repurposing the results from Algorithm 1 for previously considered sample sizes to estimate operating characteristics for new $n$ values. This process would allow us to explore new $(n,\boldsymbol{\gamma})$ combinations without conducting additional simulations. We begin the development of such a method for the design of A/B tests in Section 3. But first, we compare the FDR control provided by our design framework with the FDR control prompted by standard methods.
2.2 False Discovery Rate Control
To compare and contrast how the FDR is controlled in our design framework and under
alternative approaches, we let $V$ and $S$ be random variables that denote the number of true and false discoveries, respectively. Our design method controls the FDR such that
$\mathbb{E}_{\boldsymbol{\eta}\sim\Psi}\left\{\lim_{n\rightarrow\infty}\mathbb{E}\left[\dfrac{V}{V+S}\right]\right\}\leq q,$ (2)
where $0\div 0$ is defined to be 0. We emphasize that the FDR is controlled in expectation under the assumption that $\boldsymbol{\eta}\sim\Psi$, and we discuss the implications when this assumption does not hold true in Section 5. Our FDR control is also defined with respect to the model $f(\boldsymbol{y};\boldsymbol{\eta})$ for the data that practitioners must select. The criterion in (2) holds true as $n\rightarrow\infty$ because our method to explore the sample size space leverages asymptotic theory. Nevertheless, our method performs well with finite $n$ since we estimate sampling distributions with simulation.
The most popular frequentist method to control the FDR was proposed by Benjamini and Hochberg (1995). Their method controls the FDR such that
$\inf_{\boldsymbol{\eta}}\lim_{n\rightarrow\infty}\mathbb{E}\left[\dfrac{V}{V+S}\right]\leq q,$ (3)
where $0\div 0$ is again defined to be 0. The criterion in (3) holds for an arbitrary value for the parameter $\boldsymbol{\eta}$ that indexes the data generation process. Standard methods for FDR control assume that $p$-values are uniformly distributed under each $H_{0,k}$. This uniformity is often justified asymptotically using the central limit theorem, so standard methods for FDR control do not require practitioners to choose a model $f(\boldsymbol{y};\boldsymbol{\eta})$ for the data. However, the FDR control in Benjamini and Hochberg (1995) is only attained for certain dependence structures, including settings where all the hypotheses are independent. Benjamini and Yekutieli (2001) developed a method that attains the criterion in (3) for an arbitrary dependence structure, but simulation-based procedures are recommended for optimal control of the FDR for a specific dependence structure (Yekutieli and Benjamini, 1999). Ultimately, our design framework requires users to make more choices about the data generation process and model for the data; the trade-off of using our methods is that the recommended sample sizes and decision thresholds are optimized when these choices are reasonable.